# A Modular and Multimodal Generative AI Framework for Urban Building Energy Data: Generating Synthetic Homes

### v1

@@ 4 Conclusion
The experiments in Section 3 demonstrate a level of confidence above previous work such as Zhang et al. [19] and Xiao and Xu [20]. These works and others either accept the output of the LLMs or use procedural evaluation techniques. Occlusion and ablation offer stronger empirical support that the image processing and labeling components of our pipeline—which are both LLM-based—perform as expected. Our results directly address and mitigate concerns about hallucination and unreliability raised in prior work [24]. It is right to be skeptical of LLMs and other ML models—but with our results, we clearly demonstrate our pipeline is not as susceptible to some of the issues Liu et al. raise.
In this work, we presented a novel way to generate data for urban energy research and other fields that use similar data. Using a multimodal pipeline, we produce this labeled data in a cost-effective and efficient way. We also performed an in depth analysis to ensure that our generative AI pipeline components were behaving as expected, taking into account the concerns of Liu.
Generative AI is a powerful tool, and we recognize the importance of this new frontier as described by Xu et al. [17] and Liu. Not only does our work demonstrate a step forward in the urban energy field, it also presents a useful, malleable tool that experts and researchers can use for a variety of tasks.
Consequently, there is much work that can utilize or expand upon this study. One promising direction is the use of this pipeline to train machine learning algorithms. With the pipeline introduced here, one could create a neural network that recommends energy efficiency retrofits based on multimodal data.
Overall, the high prices, lack of availability, or privacy concerns attached to the collection of urban energy data can be avoided by generating synthetic data. With our framework, all of these concerns are alleviated and realistic, labeled data is produced. This reduced dependence on costly or restricted data sources provides a pathway towards more accessible and reproducible research.
@@ Acknowledgments
The authors extend gratitude to Lafayette College for supporting this work by funding it through the EXCEL Scholars program and providing computational resources with the Firebird high performance computing cluster. Access to the JetStream2 system was made possible via the NSF-funded ACCESS program, facilitated by Lafayette College. We also thank to the NSF, ACCESS for supporting this resource, and to Indiana University for hosting JetStream2 as a part of the ACCESS program.
@@
- Ablation
- A test used to evaluate multimodal models by observing how a model’s output responds to each varying modality.
- AI
- Artificial Intelligence
- COP
- Coefficient of Performance
- EnergyPlus
- A versatile building energy modeling tool provided by the United States Department of Energy. Orginally published in 2002.
- GeoJSON
- Geographic JSON
- GPT
- A popular large language model by OpenAI with a wide variety of uses.
- IDF
- Input Definition File
- JSON
- JavaScript Object Notation
- JSONL
- JSON Lines
- LLaVA
- An LLM integrated with CLIP, OpenAI’s visual model. Introduced by Liu et al. in 2023.
- LLM
- Large Language Model
- ML
- Machine Learning
- Occlusion
- A test commonly used to evaluate visual machine learning models by blocking out parts of an image and comparing to a baseline.
@@ A Review of
Multi-Domain Urban Energy Modelling Data, Clean Energy and
Sustainability 2 (4) (2024) 10016–10016.
@@
- [35] A. Radford, J. W. Kim, C. Hallacy, A. Ramesh, G. Goh, S. Agarwal, G. Sastry, A. Askell, P. Mishkin, J. Clark, G. Krueger, I. Sutskever, Learning Transferable Visual Models From Natural Language Supervision (Feb. 2021). arXiv:2103.00020, doi:10.48550/arXiv.2103.00020.
@@
URL https://energyplus.net/assets/nrel_custom/pdfs/pdfs_v23.2.0/AuxiliaryPrograms.pdf - [43] S. Hooker, D. Erhan, P.-J. Kindermans, B. Kim, A Benchmark for Interpretability Methods in Deep Neural Networks (Nov. 2019). arXiv:1806.10758, doi:10.48550/arXiv.1806.10758.
- [47] Y.-H. H. Tsai, S. Bai, P. P. Liang, J. Z. Kolter, L.-P. Morency, R. Salakhutdinov, Multimodal Transformer for Unaligned Multimodal Language Sequences (Jun. 2019). arXiv:1906.00295, doi:10.48550/arXiv.1906.00295.
Appendix A Labeler Experimental Values
In this appendix, we list the detailed experimental values used to test the labeler in Section 3.2. Tables 2 and 3 give the text we used as home inspection notes. Table 4 maps our EnergyPlus ablation test onto EnergyPlus variables, and Table 5 shows the value that each index represents. Finally, Table 6 gives the default values used in the EnergyPlus ablation test.
Any arbitrary simulation can be used for the textual ablation test and any arbitrary string for the EnergyPlus ablation test since both of these are constant and do not vary throughout the experiment. We pick an arbitrary simulation and use the neutral inspection note that follows:
The home has two stories with vinyl siding and a shingled roof. Windows appear to be single-hung with no visible damage. The HVAC system is located on the first floor near the utility room. Insulation levels in the attic are unknown. Doors are wood-core with standard weather stripping.
| ID | Efficiency Level | Description |
| HVAC1 | Very Inefficient | There is an older HVAC unit installed, with signs of rust on the exterior. |
| HVAC2 | Inefficient | The HVAC system appears to be in working condition but is an older standard-efficiency model. |
| HVAC3 | Moderate | The home uses window AC units rather than a central HVAC system. |
| HVAC4 | Efficient | The HVAC system was recently replaced with a standard high-efficiency model and is expected to operate efficiently. |
| HVAC5 | Very Efficient | A state-of-the-art HVAC system with smart thermostats and variable-speed compressors was recently installed, maximizing energy efficiency. |
| ID | Efficiency Level | Description |
| INS1 | Very Inefficient | Attic insulation is minimal, with exposed joists visible throughout, causing significant heat loss. |
| INS2 | Inefficient | No signs of added insulation were observed in the basement ceiling, suggesting potential energy inefficiency. |
| INS3 | Moderate | Walls appear to be adequately insulated based on construction year, though no upgrades were observed. |
| INS4 | Efficient | Blown-in insulation is present in the attic to a depth of approximately 10 inches, providing good thermal resistance. |
| INS5 | Very Efficient | High-performance spray foam insulation was installed throughout the walls, attic, and basement, providing maximum energy efficiency. |
| Variable | Corresponding EnergyPlus Variable |
| WALLR | Wall R-value |
| ROOFR | Roof R-value |
| HVACH | HVAC Heating COP |
| HVACC | HVAC Cooling COP |
| Variable | 1 | 2 | 3 | 4 | 5 |
| WALLR | 4.0 | 7.0 | 13.0 | 20.0 | 30.0 |
| ROOFR | 10.0 | 20.0 | 30.0 | 40.0 | 50.0 |
| HVACH | 0.7 | 0.8 | 0.9 | 0.95 | 1.0 |
| HVACC | 1.0 | 2.0 | 3.0 | 3.5 | 4.0 |
| Variable | Default Value |
| Air Change Rate | 2.0 |
| HVAC Heating COP | 0.8 |
| HVAC Cooling COP | 3.0 |
| Window U Value | 2.0 |
| Wall R Value | 13.0 |
| Roof R Value | 30.0 |
| HVAC System Type | Gas Furnace |
Appendix B Occlusion Results
In this appendix, we provide detailed information and results from our occlusion and reverse occlusion tests. Firstly, when running the test we utilized identical prompts for each model. For GPT, we prompted:
@@ Appendix B Occlusion Results
As noted in Section 3.1, we also ran reverse occlusion tests. While we didn’t calculate quantitative values for these results, we attach them in Table LABEL:tab:rev_occ_key.
@@
| Figure | LABEL:fig:occ_forward_bad_1 | LABEL:fig:occ_forward_bad_2 | LABEL:fig:occ_forward_bad_3 | LABEL:fig:occ_forward_bad_4 | LABEL:fig:occ_forward_bad_5 | LABEL:fig:occ_forward_bad_6 | LABEL:fig:occ_forward_bad_7 | LABEL:fig:occ_forward_bad_8 | LABEL:fig:occ_forward_bad_9 | LABEL:fig:occ_forward_bad_10 |
@@
| Figure | LABEL:fig:occ_forward_good_1 | LABEL:fig:occ_forward_good_2 | LABEL:fig:occ_forward_good_3 | LABEL:fig:occ_forward_good_4 | LABEL:fig:occ_forward_good_5 | LABEL:fig:occ_forward_good_6 | LABEL:fig:occ_forward_good_7 | LABEL:fig:occ_forward_good_8 | LABEL:fig:occ_forward_good_9 | LABEL:fig:occ_forward_good_10 |
in 1,…,10
in 1,…,10

### v2

@@ 5 Conclusion
In this work, we presented a novel way to generate multimodal data for urban energy research and other fields that use similar data. Using a multimodal pipeline, we produce this data in a cost-effective and efficient way. We also performed an in-depth analysis to ensure that our generative AIpipeline components were behaving as expected, taking into account the concerns of [24].
Generative AIis a powerful tool, and we recognize the importance of this new frontier as described by Xu et al. [17] and Liu et al [24]. Not only does our work demonstrate a step forward in the urban energy field, it also presents a useful, malleable tool that experts and researchers can use for a variety of tasks. There is much work that can utilize or expand upon this study. One promising direction is the use of this pipeline to train MLalgorithms.
For instance, a new model of the pipeline could utilize the resulting simulation data and inspection notes to label the efficiency of particular components that affect the energy efficiency of a home. These labeled homes could be used as inputs for an MLmodel that could provide recommendations for energy efficiency retrofits. Furthermore, this model, trained with our synthetic data, could be used to predict the priorities for retrofits in real houses.
Overall, the high prices, lack of availability, or privacy concerns attached to the collection of urban energy data can be avoided by generating synthetic data. With our framework, all of these concerns are alleviated and realistic data is produced. This reduced dependence on costly or restricted data sources provides a pathway towards more accessible research.
@@ Acknowledgments
The authors extend gratitude to Lafayette College for supporting this work by funding it through the EXCEL Scholars program and providing computational resources with the Firebird high performance computing cluster. Access to the JetStream2 system was made possible via the NSF-funded ACCESS program, facilitated by Lafayette College. We also thank the NSF and ACCESS for supporting this resource, and Indiana University for hosting JetStream2 as a part of the ACCESS program.
@@ A Review of
Multi-Domain Urban Energy Modelling Data 2 (4) 10016–10016.
@@
- [35] S. Hooker, D. Erhan, P.-J. Kindermans, B. Kim, A Benchmark for Interpretability Methods in Deep Neural Networks (Nov. 2019). arXiv:1806.10758, doi:10.48550/arXiv.1806.10758.
@@
URL https://energyplus.net/assets/nrel_custom/pdfs/pdfs_v23.2.0/AuxiliaryPrograms.pdf - [42] A. Goncalves, P. Ray, B. Soper, J. Stevens, L. Coyle, A. P. Sales, Generation and evaluation of synthetic patient data, BMC Medical Research Methodology 20 (1) (2020) 108. doi:10.1186/s12874-020-00977-1.
- [43] N. Patki, R. Wedge, K. Veeramachaneni, The synthetic data vault, in: 2016 IEEE International Conference on Data Science and Advanced Analytics (DSAA), 2016, pp. 399–410. doi:10.1109/DSAA.2016.49.
[44]
E. Present, P. R. White, C. Harris, R. Adhikari, Y. Lou, L. Liu, A. Fontanini,
C. Moreno, J. Robertson, J. Maguire,
Resstock 2024.1 dataset, resStock
2024.1 dataset (2024).
doi:10.2172/2319195.
URL https://doi.org/10.2172/2319195 -
[45]
J. Reyna, A. Fontanini, E. Present, L. Liu, R. Adhikari, C. Bianchi,
J. Brossman, R. Chintala, K. Clark, C. Harris, S. Horowitz, Y. Lou,
J. Maguire, N. Merket, N. Moore, P. Munankarmi, J. Robertson, N. Sandoval,
A. Speake, K. Stenger, P. White, E. Wilson,
Resstock technical reference
documentation v3.3.0, resStock Technical Reference Documentation (2025).
doi:10.2172/2583488.
URL https://doi.org/10.2172/2583488 - [46] A. Radford, J. W. Kim, C. Hallacy, A. Ramesh, G. Goh, S. Agarwal, G. Sastry, A. Askell, P. Mishkin, J. Clark, G. Krueger, I. Sutskever, Learning Transferable Visual Models From Natural Language Supervision (Feb. 2021). arXiv:2103.00020, doi:10.48550/arXiv.2103.00020.
Appendix A Occlusion Results
In this appendix, we provide detailed information and results from our occlusion tests. Firstly, when running the test we utilized identical prompts for each model. For GPT, we prompted:
@@ Appendix A Occlusion Results
Next, we provide our hypotheses pertaining to the performance of both models. We hypothesize that LLaVA’s advantage stems from its architecture: it employs an early fusion design in which images are converted into visual tokens by CLIP [46] that are then prepended to the textual prompt [28]. These visual tokens are passed through the transformer layers of the LLMtogether with text tokens, enabling multimodal attention at every layer. It has been demonstrated that vision tokens are most critical for a model in early layers, when they fuse into text tokens [47]. A likely result of this is the reduction of the model’s sensitivity to variations in the image, which may explain its stronger performance in our experiments.
@@
| Figure | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 | 15 | 16 |
@@
| Figure | 17 | 18 | 19 | 20 | 21 | 22 | 23 | 24 | 25 | 26 |
Appendix B GeoJSON Generation Prompt
In this appendix, we provide the prompt used to generate the GeoJSON data in Section 3.3.